Circular No. 7748

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001fh IN PGC 66592 S. Benetti, G. Altavilla, E. Cappellaro, A. Pastorello, M. Turatto, S.

Desidera, and E. Giro, Osservatorio Astronomico di Padova, write: “A noisy CCD spectrum (range 380–770 nm, resolution 2.5 nm) of SN 2001fh (cf. IAUC 7744), obtained on Nov. 6.97 UT with the Asiago 1.82-m telescope (+ AFOSC) under poor atmospheric conditions, is similar to that of the peculiar type-Ia SN 1986G few days before maximum (Cristiani et al. 1992, A.Ap. 259, 63). The spectrum shows a relatively red continuum, suggesting significant reddening. This is consistent with the detection of a strong absorption at 597.0 nm (equivalent width, EW, ∼ 0.35 nm) attributed to interstellar Na I D absorption in the host galaxy (V_{rec} = 3920 km/s from the LEDA Extragalactic Database). The strongest feature of the spectrum is the Si II 635.5-nm absorption measured at ∼ 621.3 nm, from which an expansion velocity of ∼ 10500 km/s is deduced.”

T. Matheson, S. Jha, P. Challis, and R. Kirshner, Center for Astrophysics, report: “A spectrum (range 370–750 nm) of SN 2001fh, obtained by P. Berlind on Nov. 7.19 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 3894 km/s for the host galaxy, the supernova expansion velocity is ∼ 11 000 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 1 ± 2 days before maximum light. The spectrum shows strong narrow Na I D absorptions at zero velocity (0.09-nm EW) and at the velocity of the host galaxy (0.16-nm EW). The Schlegel et al. (1998, Ap.J. 500, 525) dust maps indicate a Galactic reddening of E(B − V) = 0.76. In addition, the overall shape of the spectrum declines toward the blue, implying that the supernova is heavily extinguished by dust.”

SUPERNOVA 2001fu IN MCG –03-23-11 Matheson et al. also report that a spectrum (range 370–750 nm) of SN

2001fu (cf. IAUC 7746), obtained by Berlind as above on Nov. 7.51 UT, shows it to be a type-Ia supernova a few weeks past maximum. Crosscorrelation of a spectrum of the host galaxy nucleus with absorption-line velocity templates indicates a recession velocity of 1730 ± 30 km/s.

SS 433 Corrigendum. On IAUC 7747, line 3, for 0.5 mJy read 5 mJy; on line

4, for 0.05 mK read 5 mK

2001 November 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green